OpenStudy (anonymous):

8 – (3 + b) = b – 9 ? 7 8 no solution identity

OpenStudy (anonymous):

8-(3+b) = b-9 Distrib the -1 on the (3+b) because -(3+b) is really -1(3+b) right? 8-3-b=b-9 Combine like terms 5-b=b-9 Get variables on one side and numbers on the other : add b to both sides 5=2b-9 + 9 14=2b divide by 2 7=b

OpenStudy (anonymous):

:O thank you u helped me understand what im doing thanks!!!! :) and i was just curious if it was 7 but i was guessing so glad im right :)